OpenStudy (anonymous):

x+3y=-28 y=-5x This is what I've come up with so far since x=1 i'll use that 1(5x)+3y=-28

OpenStudy (anonymous):

Timesing 1*5 gives me 5x+3y=-28

OpenStudy (anonymous):

so now whaht?

OpenStudy (ivettef365):

just substitute the y for -5x you did it to the x

OpenStudy (anonymous):

So I do 5x+3(-5)=-28?

OpenStudy (anonymous):

so 5x-15=-28?

OpenStudy (anonymous):

This doesn't look right

OpenStudy (e.mccormick):

You still need to solve it for x.

OpenStudy (anonymous):

I don't understand this question at all

OpenStudy (jdoe0001):

hmmm x+3y=-28 y=-5x is that the original system of equations?

OpenStudy (e.mccormick):

Also, it looks like you subbed wrong... \(x+3y=-28 \) \(y=-5x\) Means: \(x+3(-5x)=-28 \)

OpenStudy (e.mccormick):

Solve that for x. Once you have a number, put that number into either equation in place of x and then you can solve for y.

OpenStudy (anonymous):

OK That makes somewhat more sense That would give me x-15x=-28

OpenStudy (anonymous):

Wait is that right?

OpenStudy (e.mccormick):

Yes. But it is not complete. What is x-15x?

OpenStudy (anonymous):

Um...This is easy i should know this

OpenStudy (anonymous):

x=1 right?

OpenStudy (e.mccormick):

No. Think aout this for a moment: 1-15=?

OpenStudy (anonymous):

If it is then it's -14x

OpenStudy (e.mccormick):

x means 1x, which is probably what you meant.

OpenStudy (anonymous):

Right that's what I meant :)

OpenStudy (e.mccormick):

So we are to: -14x=-28 Next step?

OpenStudy (anonymous):

x-15x=-28 so -14x=-28 -28/-14 x=2

OpenStudy (anonymous):

Then I plug in 2 into y=-5x y=-5(2) y=-10

OpenStudy (e.mccormick):

YES! Now sub in the 2 for x in one of the originals... Yah, like that. That is it.

OpenStudy (anonymous):

Thanks @e.mccormick :D

OpenStudy (e.mccormick):

np. have fun!